OpenStudy (anonymous):

ffind the slope of the line that passes through the pair of points (17,9) and (5,29)

OpenStudy (anonymous):

slope = (y2 - y1) / (x2 - x1) .....set 1 (17,9) : set 2 (5,29) slope = (29 - 9)/(5 - 17) slope = 20/- 12 slope = - 5/3 or -1.667

OpenStudy (anonymous):